# Visual Document Understanding and Question Answering: A Multi-Agent Collaboration Framework with Test-Time Scaling

### v1

@@ Abstract
Existing vision-language models (VLMs), whether generalists or specialists, remain constrained by their parameter scale, lack robust self-correction capabilities, and underperform in tasks involving long visual contexts and complex reasoning, resulting in suboptimal performance on document-based tasks. To address this, we propose MACT, a Multi-Agent Collaboration framework with Test-Time scaling, tailored for visual document understanding and visual question answering (VQA). It comprises four distinct small-scale agents, i.e., planning, execution, judgment, and answer agents, with clearly defined roles and effective collaboration. Notably, the judgment agent exclusively verifies correctness and redirects to prior agents for revisions, outperforming conventional correction strategies. To further expand the capability boundaries of the framework, we propose mixed reward modeling that balances agent-specific abilities and global collaboration, as well as agent-wise hybrid test-time scaling, which customizes different scaling strategies for each agent based on their functions. Evaluated on benchmarks spanning both document-based and non-document-based settings, our MACT shows superior performance with a smaller parameter scale without sacrificing the ability of general and mathematical tasks. Especially, it stands out in benchmarks involving long visual contexts and complicated reasoning. The three variants of MACT consistently hold the top three positions in average scores, leading in 13 of the 15 benchmarks. Code will be available at: https://github.com/YU-deep/MACT.git.
Introduction
For humans, acquiring and understanding visual information from documents—such as texts, webpages, charts, and tables—is a common and indispensable part of daily life. As a derivative task, document-based VQA provides an assessment of the understanding of both textual and visual document information as well as their relations, whose performance is quantifiable. Therefore, numerous studies aim to propose benchmarks for more complex document-based tasks of diverse types to evaluate model performance fairly and comprehensively. Specifically, the types have expanded from single-page documents, simple charts, and tables to real-world multi-page documents, complicated webpages, bar charts, pie charts, line charts, composite tables, matrix tables, and mathematical reasoning. Additionally, the form of the questions has also transformed from straightforward Yes/No, single-choice or direct information extraction questions to diversified question types, including but not limited to visual details detection, cross-page or multi-hop information extraction and summarization, comparison, multi-step calculation, intersection finding and unanswerable questions, which are answered based on the layout, structure, visual and textual contents. Thus, the latest flagship VLMs of various series are dedicated to expanding these boundaries of understanding documents.
(a) Qwen2.5-VL Series Based
(b) MiMo-VL Series Based
(c) InternVL-3 Series Based
@@ Introduction
Figure 1: Comparisons among the three variants of our MACT, their base models, and larger models in the same series.
Recently, the latest released general VLMs have showcased exceptional understanding and generalization capabilities. For instance, closed-source models, such as GPT-4o (Hurst et al. 2024), Gemini-2.0-Pro (DeepMind 2025) and Claude-3.7-Sonnet (Anthropic 2024) have surpassed human-level scores in certain tasks, while open-source models, e.g., Llama-3.2-Vision (Grattafiori et al. 2024), Qwen2.5-VL (Bai et al. 2025), MiMo-VL (Xia et al. 2025), InternVL-3 (Zhu et al. 2025) and Ovis2 (Lu et al. 2024b) have also achieved great performance. However, their performances in document-based tasks remain suboptimal due to the biases or deficiencies in the training data. To further enhance the ability, some researchers have endeavored to develop specialized VLMs. For example, UReader (Ye et al. 2023), TextMonkey (Liu et al. 2024b) and mPLUG-DocOwl2 (Hu et al. 2024b) focus on document understanding. However, their overall performance lags far behind that of the latest general VLMs with larger parameters.
To summarize, currently existing VLMs, including both generalists and specialists, exhibit three key limitations in understanding visual documents and answering questions: (1) Constrained by Parameter Scale: There are substantial performance gaps between larger-parameter VLMs and smaller counterparts in the document domain, and the potential of smaller-parameter VLMs remains underactivated. (2) Deficiency in Self-Correction Capacity: As self-correction is indispensable when humans tackle tasks involving challenging and complex documents, existing methods either fail to adequately address this capability or adopt suboptimal designs. (3) Underperformance on Long Visual Contexts and Reasoning: Results on some document-based benchmarks, which feature long visual contexts or require dense reasoning, are unsatisfactory, with notably low accuracy.
Given these issues, we propose a multi-agent collaboration framework with a dedicated test-time scaling strategy for visual document understanding and question answering. Within this framework, we employ four relatively lightweight collaborative agents: planning, execution, judgment, and answer agents. Through their collaboration and mixed reward modeling, each agent directs attention to both agent-tailored and global objectives, thereby unleashing the potential of the agents with small-scale parameters. For more efficient self-correction capacity, we design a mechanism that employs an additional judgment agent. It only checks the correctness of the generated execution plans and execution processes, with corrections implemented by the preceding agents. Additionally, we introduce an agent-wise hybrid test-time scaling approach that considers the distinct functions and features of each agent, markedly improving abilities in handling long visual contexts and complex reasoning. We train three variants of MACT based on different groups of base models using a two-stage pipeline, and evaluate them on 15 benchmarks, comprising 10 document-based and 5 non-document-based ones, to comprehensively assess document understanding as well as general and reasoning performance. Compared to models of comparable scale and larger-scale ones, the results indicate that the three variants of MACT show an average increase of at least 3.2% to 5.6%, securing the top three positions and achieving the best performance in 13 out of 15 benchmarks. Furthermore, our models have distinct advantages in long visual contexts and reasoning tasks, as further validated by results on selected non-document benchmarks.
Methodology
@@
This section provides a detailed elaboration of our proposed MACT, divided into three subsections. We first formally define the task, then introduce the four agent components: the planning agent, the execution agent, the judgment agent, and the answer agent, as well as the collaboration mechanisms among them. Finally, we detail the mixed reward modeling and agent-wise hybrid test-time scaling tailored to our multi-agent collaboration framework.
Task Definition
For the question answering task defined in this work, given a question $\mathcal{Q}$ and the its corresponding visual inputs $\mathcal{V}=\{v_{1},v_{2},\dots,v_{n}\}$, the target is to generate a correct answer output $\mathcal{O}$. This answer should leverage the information within $\mathcal{V}$ and answer the proposed question comprehensively and accurately.
@@
Figure 2: The overview of MACT. The upper part demonstrates our proposed multi-agent framework, which has four collaborative agents for tailored tasks. When the judgment agent detects mistakes, it redirects to previous agents for corrections. While the lower part illustrates the mixed reward modeling and agent-wise hybrid scaling for the multi-agent framework.
Multi-Agent Collaboration Framework
@@
As shown in the upper part of Figure 2, MACT consists of four collaborative agents: sequentially planning agent ($\mathcal{A}_{plan}$), execution agent ($\mathcal{A}_{exe}$), judgment agent ($\mathcal{A}_{judg}$), and answer agent ($\mathcal{A}_{ans}$). Each agent is role-specialized, which accomplishes its functionality and then passes the generated contents to the subsequent agent or outputs the final answers.
[Uncaptioned image] Planning Agent.
$\mathcal{A}_{plan}$ mainly focuses on the analysis and decomposition of the original question, making high-level executive plans. Largely following the analogical prompting principles (Yasunaga et al. 2024), it first generates $\mathit{N_{p}}$ relevant sample problems along with their corresponding plans simultaneously. By referring to each sample problem and its corresponding plan, it produces distinct execution plans for the current problem respectively, providing diverse pathways to accomplishment. It should be noted that this agent only yields high-level execution plans without implementation details. It ensures that $\mathcal{A}_{plan}$ can better formulate the plan from a holistic perspective, and avoid interfering with the execution of $\mathcal{A}_{exe}$. Moreover, it avoids premature binding to specific tools, enabling $\mathcal{A}_{exe}$ to select optimal resources for diverse tasks dynamically.
[Uncaptioned image] Execution Agent.
@@
$\mathcal{A}_{exe}$ is designed to execute plans step by step and output the execution process using the selected tools from the tool library. Specifically, it breaks down the steps in the plan into execution units, with information for each unit populated into a template. $\mathcal{A}_{exe}$ then executes these units sequentially, deriving the intended outputs from each. Once all steps are completed, it concatenates the full output and the execution processes of all steps in sequence, passing them to subsequent agents.
[Uncaptioned image] Judgment Agent.
@@
To enhance the quality of completing challenging tasks, humans often judge and correct mistakes that arise in planning or execution—such as reversed sequences, incorrect dependencies, redundant or missing steps, mismatches between plans and processes, and disconnects between steps. Similarly, some multi-agent systems, primarily focused on code generation, incorporate mechanisms that either (a) internally correct errors within the same agent (Yang et al. 2024; Sun et al. 2024) or (b) deploy an additional agent to handle both judgment and correction (Islam, Ali, and Parvez 2024; Li et al. 2025a). However, the internal correction mechanism (a) struggles to identify most mistakes and may fall into cognitive blind spots, as generation and correction rely on the same model. In contrast, the latter approach (b) requires the agent to possess strong capabilities for both judgment and correction, thereby necessitating models with larger parameters and more complex reward modeling designs. Furthermore, using a separate agent to regenerate mistaken components may result in incompatibilities or even conflicts with existing parts, given the autonomy of the two agents. In RL, the optimization objective for both (a) and (b), as illustrated in Figure 3, is to pass verification. This can lead to strategic production of vague statements or the omission of details, resulting in corrections that appear correct superficially but are actually misleading.
@@
To address the limitations of both approaches, we design a novel judgment strategy for multi-agent systems, which introduces an additional judgment agent to separate judgment from correction, thereby fostering specialization of labor. More specifically, $\mathcal{A}_{judg}$ is primarily responsible for assessing the correctness of steps in previously generated execution plans and processes, without engaging in direct correction. If mistakes are detected in any step, $\mathcal{A}_{judg}$ identifies the specific problematic step, provides a brief description of the error, and routes the issue to the appropriate agent—either $\mathcal{A}_{plan}$ or $\mathcal{A}_{exe}$—for correction. For each plan-process pair that is already correct or has been corrected, $\mathcal{A}_{judg}$ forwards them to $\mathcal{A}_{ans}$. This strategy, by decoupling judgment from correction, introduces a neutral judge with reduced subjective bias, whose focus remains solely on judgment rather than correction. Additionally, the reward design for the judgment agent can be intuitively simplified, without evaluating correction outcomes. To prevent infinite correction loops, the maximum number of corrections $N_{c}$ is set to 3.
@@
[Uncaptioned image] Answer Agent.
Counterintuitively, $\mathcal{A}_{ans}$ incorporates both the correct execution process and incorrect segments from prior processes. This facilitates direct focus on modifications within the corrected process, thereby preventing the omission of error-prone details.
Collaboration.
Throughout the full implementation of the multi-agent collaboration framework, visual inputs and questions are first fed into $\mathcal{A}_{plan}$, whose generated plan is then executed by $\mathcal{A}_{exe}$. $\mathcal{A}_{judg}$ judges the correctness the resulting execution plan-process pairs and outputs mistake flags. For correct pairs, $\mathcal{A}_{ans}$ outputs the final answers, while the information of mistakes will be passed to the previous agent $\mathcal{A}_{plan}$ or $\mathcal{A}_{exe}$ for correction, and the procedure repeats. For clarity in illustrating the workflow, we set the number of generated relevant plans to 1 and simplify the test-time scaling designs in Algorithm 1, and these details are elaborated in the subsequent section.
@@
0: question $\mathcal{Q}$, visual inputs $\mathcal{V}=\{v_{1},v_{2},\dots,v_{n}\}$, planning agent $\mathcal{A}_{plan}$, execution agent $\mathcal{A}_{exe}$, the tool library of the former agent $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$, judgment agent $\mathcal{A}_{judg}$, answer agent $\mathcal{A}_{ans}$, maximum number of corrections $N_{c}$
@@
8:  $p_{2}\leftarrow\mathcal{PF}\left(\mathcal{Q},\mathcal{P}_{relevant},\mathcal{M}\right)$
@@
Mixed Reward modeling
To guide RL in our multi-agent system and further enhance both the specialized and collaborative capabilities of each agent, we design a mixed reward strategy that integrates agent-specific and global outcome reward signals. Given that the four agents in the framework have distinct functions and reward signal preferences, we first incorporate agent-tailored rewards for each. In particular, for $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$, whose outputs are step-by-step, we employ an effective multi-modal reward model to yield step-wise process reward signals. These step-wise rewards provide instant feedback for each step and hierarchical rewards for the two agents, boosting accuracy on challenging tasks. For $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$, we directly use a reward model to generate one reward signal for each output of each agent. Furthermore, we apply a global reward, computed based on the final selected answer along the chosen path of the four agents. It strengthens rewards for correct paths, guiding the model to explore valid paths while mitigating adverse effects from erroneous ones. Besides, it mitigates the selfishness of agents to some extent, as individual agents tend to maximize their own rewards without regard for the global objective, and this global reward helps to alleviate such behavior.
Agent-Wise Hybrid Test-Time Scaling

### v2

# Visual Document Understanding and Reasoning: A Multi-Agent Collaboration Framework with Agent-Wise Adaptive Test-Time Scaling

@@ Abstract
The dominant paradigm of monolithic scaling in Vision-Language Models (VLMs) is failing for understanding and reasoning in documents, yielding diminishing returns as it struggles with the inherent need of this domain for document-based procedural reasoning, cognitive complexity, and factual accuracy. To this end, we introduce MACT, a Multi-Agent Collaboration framework with agent-wise adaptive Test-time scaling that pioneers a paradigm shift to procedural scaling, adapting dynamically to the functional entities of visual documents understanding and reasoning. MACT decomposes the visual document processing flow into four specialized agents, i.e., planning, execution, judgment, and answer, to resolve cognitive overload and introduce a critical self-correction loop for factual grounding. This collaborative architecture is amplified by an agent-wise adaptive test-time scaling strategy that intelligently allocates computational resources based on the complexity and redundancy of each functionality. Evaluated on multiple visual document understanding benchmarks, MACT achieves superior performance with a smaller parameter scale, adapting effectively to various document scenarios without compromising its general or mathematical reasoning capabilities. The three variants of MACT consistently attain top-three average performance rankings, with average performance enhancements of 9.9–11.5% over the base models. The source code will be released publicly.
1 Introduction
For humans, acquiring and understanding visual information from documents is an indispensable part of daily life. The corresponding tasks of visual documents, e.g., Visual Question Answering (VQA), serves as a quantifiable benchmark for a model’s ability to comprehend complex layouts, dense text, and structured data. To this end, the field has seen a rapid evolution of benchmarks, moving from simple documents to multi-page reports, complex charts, and webpages. The questions have similarly evolved from simple lookups to requiring semantic understanding and multi-hop reasoning, pushing the capability boundaries of VLMs.
(a) Comparisons on Qwen2.5-VL Series [qwen2025qwen25technicalreport]
(b) Comparisons on MiMo-VL Series [xia2025mimo]
(c) Comparisons on InternVL-3 Series [zhu2025internvl3]
@@ 1 Introduction
Figure 1: Comparisons among the three variants of MACT, the base models of these variants, and larger-scale models within the same model family, indicating the superiority of our framework over monolithic scaling paradigms.
The dominant paradigm for advancing VLM capabilities, both for documents and other general vision data, has been monolithic scaling, i.e., increasing parameter size, leveraging more high-quality training data. The approach has led to remarkable success, with phenomenal models like GPT-4o [hurst2024gpt], Gemini-2.0-Pro [gemini2024deepmind] and Claude-3.7-Sonnet [anthropic2024claude], demonstrating powerful general-purpose abilities. However, for document-based VQA, this strategy of uniform scaling is showing diminishing returns. As shown in Fig. 1, only limited performance gain is attained along parameter scaling for open-sourced models [qwen2025qwen25technicalreport, zhu2025internvl3] with exponential growth in computational overhead, suggesting that simply increasing parameter count is a brute-force solution that fails to address the core challenges of the domain.
The core problem is that documents are not monolithic entities that are suitable for uniform scaling. Unlike natural images, they possess a unique set of characteristics that make naive scaling less effective: (1) Suboptimal procedural reasoning. Document analysis is not a single action of perception but rather an inherently procedural, multi-step workflow [kutsyuruba2023document]: decomposing the question, forming a strategy, locating relevant information, and synthesizing a final answer. Monolithic models, with their single forward-pass architecture, attempt to solve this entire procedure at once, leading to less robust reasoning paths. (2) Cognitive overload. The procedural steps of document analysis demand a diverse set of specialized skills, such as global layout parsing, fine-grained text extraction, logical inference, and numerical calculation. A monolithic model is forced to encode all these disparate functions into a single set of weights, leading to cognitive overload and task interference. This leads to a compromise where the model achieves broad generalization at the expense of deep expertise in any single task. (3) Vulnerability to factual errors. The semantic meaning in documents is extremely sensitive to minor procedural mistakes [fitzgerald2012documents], even paragraph truncation errors or misaligning a table row can invalidate the entire answer. Monolithic, feed-forward models lack an internal verification or self-correction loop. An initial extraction error, however small, inevitably cascades through the reasoning process without being challenged, identified, and corrected.
These challenges reveal the need for a new paradigm: a shift from monolithic scaling to procedural scaling, where the problem is decomposed and specialize test-time scalings are applied to each entity. To achieve this, we propose MACT, a Multi-Agent Collaboration framework with agent-wise adaptive Test-time scaling, enabling a more effective scaling framework. Specifically, it deconstructs the monolithic model into four specialized agents that explicitly mirror the required cognitive functions. Within this framework, we employ four relatively collaborative agents: planning, execution, judgment, and answer agents. This design inherently enforces procedural reasoning, dedicating a specialist to each cognitive step from planning to synthesis. This decomposition is what resolves the cognitive overload problem, allowing each agent to achieve mastery over its specific function. The key to unlocking this specialized performance is our agent-wise adaptive test-time scaling approach. Rather than uniform scaling, it is intelligent, on-demand allocation of computational resources, allowing the execution agent to apply maximum scrutiny to a single data point while the planning agent operates efficiently on the whole document. Finally, the framework directly confronts the high cost of errors. The judgment agent acts as a built-in verification, providing the self-correction capabilities that monolithic models lack.
We train three variants of MACT based on different groups of base models using a two-stage pipeline, and evaluate them on 15 benchmarks, comprising 10 document-based and 5 non-document-based ones. Our MACT showcases superior document understanding and reasoning without compromising general capabilities. Compared to models of comparable scale and larger-scale ones, the three variants of MACT demonstrate an average increase of 3.2% to 5.6%, securing the top three positions and achieving the best performance in 13 out of 15 benchmarks. The results demonstrate that monolithic scaling is not the optimal approach for visual document-based tasks. In contrast, the adaptive test-time scaling in multi-agent framework emerges as a more advantageous alternative, as it achieves superior scaling performance.
2 Methodology
@@
This section provides a detailed elaboration of our proposed MACT, divided into three subsections. We first introduce the multi-agent framework, which decomposes the document-tailored workflow into four functionalities: planning, execution, judgment, and answer, and defines the collaboration mechanisms among them. Then, we detail the agent-wise adaptive test-time scaling and mixed reward modeling tailored to the multi-agent collaboration framework. More details are presented in Appendix 7.
@@
Figure 2: The overview of MACT. The upper part demonstrates our procedural framework, with four tailored and collaborative agents to conduct the process of document analysis. When the judgment agent detects mistakes, it redirects to previous agents for corrections. The lower part illustrates the agent-wise adaptive test-time scaling and mixed reward modeling for the multi-agent framework.
2.1 Multi-Agent Collaboration Framework
@@
Document understanding and reasoning constitutes not merely a unilateral perceptual activity, but an expected multistage process [kutsyuruba2023document]: decomposing the question, formulating execution strategies, locating information pertinent to the document, conducting validation with necessary corrections, and synthesizing a final response. As shown in the upper part of Fig. 2, MACT consists of four collaborative agents: sequentially planning agent ($\mathcal{A}_{plan}$), execution agent ($\mathcal{A}_{exe}$), judgment agent ($\mathcal{A}_{judg}$), and answer agent ($\mathcal{A}_{ans}$). Each agent is role-specialized, which accomplishes its functionality and then passes to the subsequent agent or outputs the final answers.
[Uncaptioned image] Planning Agent. Given a question $Q$ and corresponding visual document inputs $\mathcal{D}$, $\mathcal{A}_{plan}$ mainly focuses on the analysis and decomposition of the original question, making high-level executive plans on the document-based tasks. Before generating the plans for the original task, we generate $\mathit{N_{p}}$ similar and relevant instances, and their plans simultaneously, which is adapted from analogical prompting principles [michihiro2024large]. For each generated example and corresponding plan, $\mathcal{A}_{plan}$ generates a distinct execution plan respectively, which could be formulated as:
$\mathcal{P}_{rel}$ $=\mathcal{A}_{\text{plan}}\left(\mathcal{Q},\mathcal{D}\right),$ (1)
$\mathcal{P}$ $=\mathcal{A}_{\text{plan}}\left(\mathcal{Q},\mathcal{P}_{rel},\mathcal{D},\mathcal{M}\right),$ (2)
where generated relevant plans $\mathcal{P}_{rel}=\left\{r_{1},r_{2},\dots,r_{\mathit{N_{p}}}\right\}$, and execution plans $\mathcal{P}=\left\{p_{1},p_{2},\dots,p_{\mathit{N_{p}}}\right\}$, which provide different pathways to accomplishment. Besides, $\mathcal{M}$ is the mistake from $\mathcal{A}_{judg}$ for correction, and it is initialized to empty. For each relevant plan and execution plan $p$, it composed of steps $\left\{s_{1},s_{2},\dots,s_{n}\right\}$. It should be noted that we only yield high-level execution plans with limited details in this agent. To be more precise, the high-level plan describes the expected targets and requirements of each step, but does not output specific execution processes directly. It ensures that $\mathcal{A}_{plan}$ can better project the overall plan instead of execution details and would not interfere with the execution of $\mathcal{A}_{exe}$. Besides, it avoids premature commitment to specific tools, enabling $\mathcal{A}_{exe}$ to dynamically select optimal resources for diverse tasks.
@@
[Uncaptioned image] Execution Agent. $\mathcal{A}_{exe}$ is designed to execute plans step by step and output the execution process using the selected tools from the tool library $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$ (see in Appendix 7.2). Specifically, it breaks down the execution plan and each step will be regarded as an execution unit, with information for each unit populated into a template and comprised of: (1) a specific definition; (2) expected target and output; (3) existing inputs or results from the previous step. $\mathcal{A}_{exe}$ then executes the units sequentially, deriving the intended outputs from each. The execution of single step could be presented as:
$e_{i}=\mathcal{A}_{\text{exe}}\!\left(\mathcal{Q},\mathcal{D},s_{i},\mathcal{T},\mathcal{M}\right).$ (3)
Once all steps are completed, it concatenates the full output and the execution processes of all steps in sequence, passing $\mathcal{E}=\left\{e_{1},e_{2},\dots,e_{n}\right\}$ to subsequent agents.
@@
[Uncaptioned image] Judgment Agent. Document-based visual tasks exhibit high sensitivity to process-related errors [fitzgerald2012documents], especially in monolithic models that omit verification and correction procedures. Trivial procedural discrepancies inevitably distort the overall reasoning trajectory, leading to a snowball effect that results in wholly erroneous outputs. Thus, some multi-agent systems incorporate mechanisms that either (a) internally correct errors within the same agent [yang2024matplotagent, sun2024pearl] or (b) deploy an additional agent to handle both judgment and correction [islam2024mapcoder, li2025metal]. However, the (a) internal correction mechanism struggles to identify most mistakes and may fall into cognitive blind spots, as generation and correction rely on the same model. In contrast, the latter approach (b) requires the agent to possess strong capabilities for both judgment and correction, thereby necessitating models with larger parameters and more complex reward modeling designs. Furthermore, using a separate agent to regenerate mistaken components may result in incompatibilities or even conflicts with existing parts, given the autonomy of the two agents. In RL, the optimization objective for both (a) and (b), as illustrated in Fig. 3, is to pass verification. This can lead to strategic production of vague statements or the omission of details, resulting in corrections that appear correct superficially but are actually misleading. And utilizing another agent to regenerate the mistaken part might be incompatible or even causing conflicts with the existing parts because of the independence of two autonomous agents.
@@
To address the limitations of both approaches, we design a simple but effective judgment strategy for multi-agent systems, which introduces an additional judgment agent to separate judgment from correction, thereby fostering specialization of labor. More specifically, $\mathcal{A}_{judg}$ is primarily responsible for assessing the correctness of steps in previously generated execution plans and processes, without engaging in direct correction. The judgment process is summarized as:
$\mathcal{J}=\mathcal{A}_{judg}\left(\mathcal{Q},\mathcal{P},\mathcal{E}\right),$ (4)
where $\mathcal{J}=\left\{flag_{plan},flag_{exe},\mathcal{M}\right\}$, $flag_{plan}=flag_{exe}\in\left\{true,false\right\}$, and $\mathcal{M}$ is the mistake description. If mistakes are detected in any step, $\mathcal{A}_{judg}$ identifies the specific problematic step, provides a brief description of the mistake, and routes the issue to the appropriate agent, either $\mathcal{A}_{plan}$ or $\mathcal{A}_{exe}$, for correction.
For each plan-process pair that is already correct or has been corrected, $\mathcal{A}_{judg}$ forwards them to $\mathcal{A}_{ans}$. This strategy, by decoupling judgment from correction, introduces a neutral judge with reduced subjective bias, whose focus remains solely on judgment rather than correction. Additionally, the reward design for the judgment agent can be intuitively simplified, without evaluating correction outcomes. To prevent infinite correction loops, the maximum number of corrections $N_{c}$ is set to 3.
@@
[Uncaptioned image] Answer Agent. The primary function of $\mathcal{A}_{ans}$ is to output the final answer through both the incorrect processes and the correct process. Counterintuitively, $\mathcal{A}_{ans}$ incorporates both the correct execution process and incorrect segments from prior processes. This facilitates direct focus on modifications within the corrected process, thereby preventing the omission of error-prone details. Additionally, the mistake-correction pair presents a natural and complete logical closed-loop structure, which benefits answer generation based on the execution process. Thus, the final output answer is:
$\mathcal{O}=\mathcal{A}_{ans}\left(\mathcal{Q},\mathcal{E},\mathcal{M}\right).$ (5)
Collaboration. Throughout the full implementation of the multi-agent collaboration framework, visual inputs and questions are first fed into $\mathcal{A}_{plan}$, whose generated plan is then executed by $\mathcal{A}_{exe}$. $\mathcal{A}_{judg}$ judges the correctness the resulting execution plan-process pairs and outputs mistake flags. For correct pairs, $\mathcal{A}_{ans}$ outputs the final answers, while the information of mistakes will be passed to the previous agent $\mathcal{A}_{plan}$ or $\mathcal{A}_{exe}$ for correction, and the procedure repeats. For clarity in illustrating the workflow, we set the number of generated relevant plans to 1 and simplify the test-time scaling designs in Algorithm 1, and these details are elaborated in the subsequent section.
@@
0: question $\mathcal{Q}$, visual document inputs $\mathcal{D}=\{v_{1},v_{2},\dots,v_{n}\}$, planning agent $\mathcal{A}_{plan}$, execution agent $\mathcal{A}_{exe}$, the tool library of the former agent $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$, judgment agent $\mathcal{A}_{judg}$, answer agent $\mathcal{A}_{ans}$, maximum number of corrections $N_{c}$
@@
8:  $p_{2}\leftarrow\mathcal{PF}\left(\mathcal{Q},\mathcal{P}_{rel},\mathcal{M}\right)$
@@
2.2 Agent-Wise Adaptive Test-Time Scaling